OpenStudy (anonymous):

solve for x in terms of a,b, and c: ax - 3b =c 1: a(c+3b) 2: a (c-3b) 3: c-3/a 4: c+3b/a

jimthompson5910 (jim_thompson5910):

Add 3 to both sides, then divide both sides by 'a' to isolate x \[\Large ax - 3b = c\] \[\Large ax = c+3b\] \[\Large x = \frac{c+3b}{a}\]

jimthompson5910 (jim_thompson5910):

*Add 3b to both sides* (not 3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i add 3b to what?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

to c and just leave it like c+3b like I showed

OpenStudy (anonymous):

ok.....

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

then what?

jimthompson5910 (jim_thompson5910):

It's already done, look at my first post

OpenStudy (anonymous):

so i divide both sides by a?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes in that last step

OpenStudy (anonymous):

and thats how you got c+3b/a?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

so basically, collect the terms in x on one side of the equation, and the literal terms on the other side @jim_thompson5910

jimthompson5910 (jim_thompson5910):

bingo, you want x on its own side (preferably the left) and everything else on the other side

OpenStudy (anonymous):

got it, Thanks :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw